# Resolution of Erdős Problem #728: a writeup of Aristotle’s Lean proof

### v1

@@ Abstract
We provide a writeup of a resolution of Erdős Problem #728; this is the first Erdos problem (a problem proposed by Paul Erdős which has been collected in the Erdos Problems website [2]) regarded as fully resolved autonomously by an AI system. The system in question is a combination of GPT-5.2 Pro by OpenAI and Aristotle by Harmonic, operated by Kevin Barreto. The final result of the system is a formal proof written in Lean, which we translate to informal mathematics in the present writeup for wider accessibility.
The proved result is as follows. We show a logarithmic-gap phenomenon regarding factorial divisibility: For any constants $0<C_{1}<C_{2}$ there exist infinitely many triples $(a,b,n)\in\mathbb{N}^{3}$ such that
@@ 1 Introduction
Fix $0<C_{1}<C_{2}$. There exist infinitely many $(a,b,n)\in\mathbb{N}^{3}$ such that
@@ 6 Correspondence with the Lean development
While the mathematical and literature contents of the writeup is already complete, this proof has a significance of being the first recognized Erdos problem (a problem proposed by Paul Erdos and collected on the Erdos problems website https://www.erdosproblems.com) that was solved autonomously by an AI system, with no prior literature found as of the date of this writing.
As this is a milestone, it is of interest to ask about the nature of the proof (e.g. how complex/novel it is compared to other mathematical results). Unfortunately (or fortunately) the resulting proof is in formalized form in Lean. While this boosts confidence in correctness, it may not be readily digestible.
@@
It is of interest to record the story of this proof as follows. The “raw” story may be read in the Erdos problems website thread: https://www.erdosproblems.com/728.
On Jan 4, 2026, Kevin Barreto announced that he has a proof from the AI system: Aristotle by Harmonic. This is a Lean system, and the input to Aristotle is based on an informal argument from GPT-5.2 Pro. However there is one nuance here: at that time, the problem as stated on the website is vague, with the intended interpretation from Erdos unclear. Barreto’s version resolves one version of the problem, but it wasn’t clear whether this is the intended version.
Afterwards, forum participants commented on this vagueness; a consensus then emerged that this should be regarded as a partial result, which does not yet fully resolve the problem.
But - on Jan 5, 2026, Barreto again asked GPT-5.2 Pro whether its argument can be upgraded to tackle the version that forum participants had now identified as the problem. GPT-5.2 Pro responded in affirmative; Aristotle was then run based on this latest response and produced a formal proof on Jan 6, 2026.
There was a misunderstanding around this time that this second output had been based on a human’s mathematical observation, which would render the result not autonomous by the AIs involved. Nevertheless, Barreto has clarified that this was a misunderstanding and a human observation was not provided in order to reach this result. With autonomy confirmed, the problem was now understood to be fully resolved by AI.
@@
Afterwards, a forum participant who goes by KoishiChan attempted a literature review to locate any prior literature resolving this problem. KoishiChan has been highly successful in locating prior literature in the past for problems claimed to be solved by AI; however this time around an existing literature resolving this problem was not found. It is of course possible that such literature exists, as it is impossible to disprove the possibility, but we have not found such literature as of the date of this writing. Thus, the current status is that Erdos Problem 728 was fully resolved autonomously by AI with no prior human literature resolving the problem found.
@@
For reference, the conversation with ChatGPT that the author collaborates with in producing this writeup can be accessed here [^2] https://chatgpt.com/share/696148ea-ffd4-8010-830f-7c04f3ab5cb9. Note that while there are many drafts there, none of them is the current writeup, in which the author further applies significant rewrites in various places by himself.

### v2

@@ Abstract
We provide a writeup of a resolution of Erdős Problem #728; this is the first Erdős problem (a problem proposed by Paul Erdős which has been collected in the Erdős Problems website [2]) regarded as fully resolved autonomously by an AI system. The system in question is a combination of GPT-5.2 Pro by OpenAI and Aristotle by Harmonic, operated by Kevin Barreto. The final result of the system is a formal proof written in Lean, which we translate to informal mathematics in the present writeup for wider accessibility.
The proved result is as follows. We show a logarithmic-gap phenomenon regarding factorial divisibility: For any constants $0<C_{1}<C_{2}$ and $0<\varepsilon<1/2$ there exist infinitely many triples $(a,b,n)\in\mathbb{N}^{3}$ with $\varepsilon n\leq a,b\leq(1-\varepsilon)n$ such that
@@ 1 Introduction
Figure 1: Plots showing the required inequality $v_{p}(\binom{m+k}{k})\leq\kappa_{p}(m)$; we want to find places where the red is below the blue. Both plots have $m\in[1000,2000]$ and $k=10$. The two plots are for primes $p=2$ and $p=13$, respectively. Note that these plots are smoothed by a simple moving average with window 2525, since the unsmoothed plots are dense and difficult to read. Plotting code was partially provided by ChatGPT.
@@
Fix $0<C_{1}<C_{2}$. Let $0<\varepsilon<1/2$. There exist infinitely many $(a,b,n)\in\mathbb{N}^{3}$ with $\varepsilon n\leq a,b\leq(1-\varepsilon)n$ such that
@@ 6 Correspondence with the Lean development
While the mathematical and literature contents of the writeup is already complete, this proof has a significance of being the first recognized Erdős problem (a problem proposed by Paul Erdős and collected on the Erdős problems website https://www.erdosproblems.com) that was solved autonomously by an AI system, with no prior literature found as of the date of this writing.
As this is a milestone, it is of interest to ask about the nature of the proof, e.g. how complex/novel it is compared to other mathematical results. Unfortunately (or fortunately) the resulting proof is in formalized form in Lean. While this boosts confidence in correctness, it may not be readily digestible.
@@
It is of interest to record the story of this proof as follows. The “raw” story may be read in the Erdős problems website thread: https://www.erdosproblems.com/728.
On Jan 4, 2026, Kevin Barreto announced that he had a proof from the AI system: Aristotle by Harmonic. This is a Lean system, and the input to Aristotle was based on an informal argument from GPT-5.2 Pro. However there was one nuance here: at that time, the problem as stated on the website was vague, with the intended interpretation from Erdős unclear. Barreto’s version resolved one version of the problem, but it wasn’t clear whether this was the intended version.
Afterwards, forum participants commented on this vagueness; a consensus then emerged that this should be regarded as a partial result, which did not yet fully resolve the problem.
But - on Jan 5, 2026, Barreto again asked GPT-5.2 Pro whether its argument could be upgraded to tackle the version that forum participants had now identified as the problem. GPT-5.2 Pro responded in affirmative; Aristotle was then run based on this latest response and produced a formal proof on Jan 6, 2026.
There was a misunderstanding around this time that this second output had been based on a human’s mathematical observation, which would render the result not autonomous by the AIs involved. Nevertheless, Barreto had clarified that this was a misunderstanding and a human observation was not provided in order to reach this result. With autonomy confirmed, the problem was now understood to be fully resolved by AI.
@@
Afterwards, a forum participant who goes by KoishiChan attempted a literature review to locate any prior literature resolving this problem. KoishiChan has been highly successful in locating prior literature in the past for problems claimed to be solved by AI; however this time around an existing literature resolving this problem was not found. It is of course possible that such literature exists, as it is impossible to disprove the possibility, but we have not found such literature as of the date of this writing. Thus, the current status is that Erdős Problem 728 was fully resolved autonomously by AI with no prior human literature resolving the problem found.
@@
For reference, the conversation with ChatGPT that the author collaborated with in producing this writeup can be accessed here [^2] https://chatgpt.com/share/696148ea-ffd4-8010-830f-7c04f3ab5cb9. Note that while there are many drafts there, none of them is the current writeup, in which the author further applied significant rewrites in various places by himself.